OpenStudy (anonymous):

Suppose that a sphere with radius 5a has the same volume as a cone of radius 3a. What is the height of the cone? Give your answer in terms of a. Use the formulas and for the volumes of a sphere and cone respectively.

OpenStudy (anonymous):

heres the formulas sphere=v=3/3pier^3 cone=v=1/3pier^2

OpenStudy (paxpolaris):

\[\Large \frac 43 \pi \left( 5a \right)^3=\frac 13 \pi \left( 3a \right)^2h\] solve for h

OpenStudy (anonymous):

how do i figure out how to solve it?

OpenStudy (anonymous):

use just the algebra inverse operation steps....you want the h remaining alone on the right hand side, so first I would multiply each side by the reciprocal of 1/3, that being a 3/1 .....

OpenStudy (anonymous):

write that down and follow me as I go...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that would look like...\[3/1 * 4/3\Pi ^{}\]

OpenStudy (anonymous):

darn, where's the rest I typed!! \[(5a)^{3} = \Pi (3a)^{2}h\]

OpenStudy (anonymous):

is that last symbol pie

OpenStudy (anonymous):

yes, believe it or not! not a good choice..maybe its really the Roman numeral for 2! anyway pi...:-)

OpenStudy (anonymous):

ok, so simplify and on the left, the 3's cancel out when you cross cancel in that fraction, leaving you with just a 4 pi ( )^3 = pi ( ) ^2 h

OpenStudy (anonymous):

now just divide both sides by pi, and they will completely cancel out....leaving you with jsut the 4(5a)^3 = (3a)^2 h

OpenStudy (anonymous):

Now a couple of ways to go....I'll do the long way so you can understand what I do without me skipping any steps...

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Simplify each side before you do any more inverse operations (this actually could have been your very first step)....you get 4 * 125a^3 = 9a^2*h see where this all came from?

OpenStudy (anonymous):

so far yes

OpenStudy (anonymous):

ok, now keep simplifying,,,LS will = 1000a^3 = 9a^2 h

OpenStudy (anonymous):

now divide both sides by 9^2....you get (1000a^3)/(9a^2) = h

OpenStudy (anonymous):

simplify another time, to get 1000a/9 = h.....do you have a multiple choice?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

is that one of them?? I hope! :-)

OpenStudy (anonymous):

no there is 500 on top the fraction

OpenStudy (anonymous):

but the 9's there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh! mis-multiplied! that's what I get for doing all this in my head....4* 125 is 500, not 100. :-)

OpenStudy (anonymous):

not 1000

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Have any others similar to this?

OpenStudy (anonymous):

i got 500a over 9 and 9 over 500a

OpenStudy (anonymous):

500a/9 is the correct answer......you could also simplify the ( ) 1st, then cross cancel on each side with the fractions you have, first the one on the left then the one on the right....

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

you're welcome!